### v2

# Adaptive Self-improvement LLM Agentic System for ML Library Development

Abstract
ML libraries, often written in architecture-specific programming languages (ASPLs) that target domain-specific architectures, are key to efficient ML systems. However, writing these high-performance ML libraries is challenging because it requires expert knowledge of both ML algorithms and the ASPL. Large language models (LLMs), on the other hand, have shown general coding capabilities. However, challenges remain when using LLMs for generating ML libraries using ASPLs because 1) this task is complicated even for human experts and 2) there are limited code examples due to the esoteric and evolving nature of ASPLs. We present an adaptive self-improvement agentic system that enables LLMs to perform such complex reasoning under limited data by iteratively improving their capability through self-generated experience. In order to evaluate the effectiveness of our system, we construct a benchmark of a typical ML library and generate ASPL code with both open and closed-source LLMs on this benchmark. Our results show improvements of up to $3.9\times$ over a baseline single LLM  [^1] The example code is public at https://github.com/zhang677/PCL-lite.
LLM agents, Self-improvement learning, Machine learning library
1 Introduction
Figure 1: We propose an adaptive self-improvement LLM agentic system. LLM agents start from their base knowledge and accumulate experiences through parallel sampling. Our adaptive self-improvement learning algorithm filters high-quality answers, stratifies the earned experiences by difficulty, and adaptively selects demonstrations to enhance LLM agents.
With the ending of Dennard Scaling and Moore’s Law, computer architectures are specializing in domain applications to achieve greater performance and efficiency and will continue to do so (Hennessy & Patterson, 2019). New domain-specific architectures (DSA) typically come with new architecture-specific programming languages (ASPL), such as CUDA for NVIDIA GPUs (NVIDIA, 2025), HIP for AMD GPUs (AMD, 2025), and Pallas for Google TPUs (Google, 2025). Even existing ASPLs change as generations of DSAs evolve because new DSAs introduce specialized functions to these existing ASPLs (Choquette, 2023). Efficiently utilizing these new functions requires fundamentally different programming styles and thus new ASPLs (Thakkar et al., 2023; Hagedorn et al., 2023).
Each DSA needs a corresponding ML library, a collection of ML operators written in its ASPL, before programmers can effectively use the DSA to accelerate ML applications. ML library development is challenging as it requires expertise in both ML algorithms and the target ASPL. Essentially, library development is a generation process that composes low-level ASPL primitives into high-level ML operators (Dong et al., 2024; Ye et al., 2025).
Furthermore, ML library development using ASPLs requires complex reasoning while minimizing data requirements. ML libraries development often occurs in parallel with hardware manufacturing to meet production deadlines (Villa et al., 2021). This time-constrained library–chip co-design process leaves limited code examples. Moreover, this task is complicated even for expert human programmers. For example, the publication of FlashAttention-3 (Shah et al., 2024) lagged behind the release of the H100 by two years. The challenge is further intensified by the need for ML libraries to co-evolve with new hardware to sustain performance. For instance, directly adapting FlashAttention-2 (Dao, 2023) from the A100 to H100 GPU witnessed a 47% performance drop (Spector et al., 2024).
The challenges in ML library development call for more automatic solutions. Furthermore, these automatic solutions need to self-improve to perform complex reasoning starting from simple and limited data. Large language models (LLMs) have demonstrated emerging capabilities in code generation (Kaplan et al., 2020; Wei et al., 2022a). Moreover, empirical evidence implies that LLMs already have the base knowledge of ML algorithms (Ouyang et al., 2024). Therefore, we explore the use of LLM agents to develop ML libraries with emerging ASPLs.
Current self-improvement methods for LLM agents fall short because of limited exploration or low data efficiency. LLM agents can enhance their performance by synthesizing semantically similar data (Yu et al., 2023; Shinn et al., 2024; Zhao et al., 2024). Although these methods are effective for local exploration (Chen et al., 2024), they are insufficient for tasks that require substantial cognitive effort (Huang et al., 2023). Self-improvement learning can significantly improve reasoning ability through reinforcement learning (Cobbe et al., 2021; Bai et al., 2022; Singh et al., 2023). This approach, however, currently requires hundreds of effective trajectories sampled from LLM agents for each problem (Wang et al., 2024), making it unsuitable for complex scenarios with limited data availability.
To address these limitations, we design an adaptive self-improvement learning algorithm integrated with an agentic system organization. This approach not only produces a self-improving agentic system to assist humans but also generates high-quality ML operators that can be leveraged by other systems. We show our system in Figure 1. Similar to human experiential learning (Kolb, 2014), our techniques create a self-improvement cycle: LLM agents evolve through earned experiences and these evolved agents can earn more experiences. This self-improvement cycle is fully automated, involving no human experts beyond the ASPL designers themselves, who initially tell the models how to use the ASPL primitives.
Our algorithm prioritizes hard-earned experiences gained from completing challenging tasks, inspired by curriculum learning (Bengio et al., 2009). When these hard-earned experiences are exhausted, the algorithm adaptively increases the number of demonstrations by incoporating experiences from less challenging tasks. Section 6.1 shows that hard-earned experiences improve LLM agents more efficiently than mixed ones. While mixed experiences may dilute demonstrations, they can help agents overcome learning obstacles and complete more tasks. As a byproduct, the algorithm adaptively increases test-time compute on challenging tasks until they are complete or the data is exhausted.
To emulate the library–chip co-design process, we choose Streaming Tensor Programs (STeP) as our target ASPL for library generation. STeP is an emerging ASPL designed for next-generation reconfigurable dataflow architectures (Prabhakar et al., 2017), a family of DSAs for AI (Prabhakar & Jairath, 2021; Chen et al., 2023; Prabhakar et al., 2024). The only public document of STeP is a non-archival three-page workshop publication (Sohn et al., 2024a), which defines its semantics without any code examples or execution environments. Therefore, STeP programs do not exist in the training corpus of any LLM.
Figure 2: Portion of completed tasks (Pass@n) across models using single LLM, agentic system, and adaptive self-improvement agentic system, highlighting performance improvement.
Putting all these together, our system solves up to 96% of the tasks in our benchmark and achieves up to a $3.9\times$ improvement over a baseline single LLM, as shown in Figure 2. The contributions of this paper are: (1) an adaptive self-improvement learning algorithm that enables LLM agents to continuously construct ML libraries through adaptive experience-driven evolution; (2) an end-to-end agentic system that uses adaptive self-improvement to develop an ML library for STeP, an ASPL for a next-generation AI accelerator; (3) a complete evaluation of the adaptive self-improvement learning algorithm and the integrated agentic system on a realistic benchmark constructed from common ML operators.
Figure 3: We benchmarked eight groups of ML operators within a common LLM layer (Jiang et al., 2024). These operators are categorized as dynamic, static matrix, and static vector, with parentheses indicating the number of tasks associated with each group.
2 Background
In this section, we provide background on how DSAs are programmed through their ASPLs, describe how these ASPLs are used to create end-user ML libraries, and identify key challenges of this library generation process. We also establish STeP as the target ASPL to explore LLM techniques for ML library development. Key concepts related to the background are listed in Table 8 for reference.
2.1 Architecture-specific programming languages
ASPLs describe the low-level programming interface of a DSA using primitives and specialized functions. Primitives model the basic execution pattern similar to general-purpose programming language constructs, and specialized functions control specialized accelerator units that are optimized for domain applications on the DSA. Unlike domain-specific languages (DSLs), which are a top-down distillation of the domain algorithms, ASPLs refer to a bottom-up abstraction of the underlying chip architecture.
2.2 ML library development using ASPLs
ML libraries developed in ASPLs face portability challenges because ASPLs rapidly evolve to align with DSA updates in new generations to meet the demands of growing ML workloads. For example, the matrix multiplication units on NVIDIA GPUs and their corresponding MMA instructions have been updated every generation of GPU since their introduction (NVIDIA, 2017). Consequently, every library function that uses MMA instructions must be rewritten in a new ASPL (e.g. CUDA with new instructions) per generation. Moreover, ML libraries need to be shipped at the same time as the chip. In this case, library development costs are no longer negligible.
To solve these challenges, we propose to enhance users’ learning capabilities for a given ASPL. This approach offers an alternative to current automation techniques that focus on simplifying the learning curve for new ASPLs. Mainstream automation techniques compromise by optimizing ML operators whose performance is most significantly affected by the ASPL update (Tillet et al., 2019; Thakkar et al., 2023). Only focusing on certain ML operators does not fully incorporate some ASPL updates, such as memory optimizations, which could potentially accelerate any ML operator. Meanwhile, new ML operators are being proposed (Gu & Dao, 2023; Sun et al., 2024). Given these factors, we need better automation to improve the productivity of ML library development using APSLs.
Figure 4: STeP is an ASPL for next-generation RDAs that models streaming dataflow execution. (a) STeP example program for a simplified MoE module. More details can be found in Section A.1. (b) An illustration of the streaming dataflow execution of (a) on an RDA of memory units (MU), and compute units (CU). (c) The SN40L, a deployed RDA chip.
2.3 STeP for next-generation RDA
We chose STeP as our target ASPL due to its potential for better efficiency and status as a research prototype ASPL. STeP’s efficiency potential stems from its role as an ASPL for next-generation RDAs, which are a promising alternative to GPUs. The SN40L, a deployed RDA implementation shown in Figure 4(c), demonstrates record-breaking inference speeds for the Llama 3.1 405B model (SambaNova, 2024). Although STeP does not yet have a path to a fabricated chip, developing ML libraries in STeP still presents similar challenges as other ASPLs. Writing STeP programs requires complex reasoning about streaming dataflow execution, and our work began without any existing executable STeP programs to reference.
Similar to other ASPLs, primitives and specialized functions compose to form operational semantics in STeP. Specifically, STeP primitives describe different stream token manipulation strategies, and STeP specialized functions express different configurations of RDA units. Inspired by parallel patterns and array programming (Hsu et al., 2023; Rucker et al., 2024), the streaming dataflow execution model in STeP treats data as streams of tokens flowing between computational units in space. STeP extends the conventional dataflow execution model of RDAs by introducing streaming semantics, as shown in Figure 4(b). This approach unifies data values and control signals into stream tokens, embedding control flow directly into the data to enable greater dynamism. A stream can be consumed by at most one primitive because of the queueing nature of dataflow (Zhang et al., 2021), which is called an affine type constraint in programming language theory (Wikipedia, 2024). The affine type constraint is a global property of the program since it counts the usage of a variable in the whole program.
STeP primitives are categorized as either arithmetic or shape manipulation. Arithmetic primitives apply computations and control flow to stream tokens. Shape manipulation primitives reshape the data within the stream by changing the control tokens.  Figure 4(a) and Figure 17 are example STeP programs that contain 5 arithmetic primitives and only shape manipulation primitives, respectively.
Algorithm 1 Adaptive self-improvement learning
[…]
3:repeat
4:  $\mathcal{E}\leftarrow\beta(\mathcal{D},m)$ $\triangleright$ stratification
5:  for $e_{j}\in\mathcal{E}$ do
6:   $d_{j}\leftarrow[e_{0},e_{1},...e_{j}]$ $\triangleright$ selection
7:   // Parallel sampling
8:   $\mathcal{C}_{t}\leftarrow\{\mathbb{E}_{y\sim p_{\theta}(y|x_{i},d_{j})}[r(x_{i},y)]\mid x_{i}\in\mathcal{X}\}$
9:   $\mathcal{B}_{t}\leftarrow\{(x_{i},y)\mid r(x_{i},y)=1,x_{i}\in\mathcal{X}\}$
10:   $\mathcal{S}_{t}\leftarrow\{x_{i}\mid c_{i}>0,c_{i}\in\mathcal{C}_{t}\}$
11:   if $\mathcal{B}_{t}\neq\emptyset$ then
12:    $\mathcal{D}\leftarrow\mathcal{D}\cup\sigma(\mathcal{B}_{t},\mathcal{C}_{t},\mathcal{S}_{t})$ $\triangleright$ filtering
13:    $\mathcal{X}\leftarrow\mathcal{X}\setminus\mathcal{S}_{t}$
14:    $t\leftarrow t+1$
15:    break
16:   end if
17:   $t\leftarrow t+1$
18:  end for
19:until $\mathcal{X}=\emptyset\lor(d_{j}=\mathcal{D}\land\mathcal{B}_{t-1}=\emptyset)$
19: Solutions: $\mathcal{D}$
Figure 5: Running example of Algorithm 1 with $m=3$ and $|\mathcal{X}|=5$. Orange circles are demonstrations $d_{j}$ for the current iteration. Green circles are finished tasks and white ones are not. $m=3$ means $\beta$ stratifies demonstrations into 3 levels: hard, medium, and easy. Iter 1 and 3 consider hard-only examples, and the other iterations consider mixed examples.
3 Adaptive self-improvement learning
Our adaptive self-improvement learning evolves LLM agentic systems with data generated by the systems themselves. This algorithm samples the agentic system in parallel for correct answers with success rates, filters high-quality correct answers, stratifies the earned experiences, and adaptively updates demonstrations until all the tasks are solved or the demonstrations are exhausted. The complete algorithm for adaptive self-improvement learning is shown in Algorithm 1. As a byproduct, the algorithm adaptively assigns more test-time compute to harder tasks by excluding tasks from the task set after completion. Figure 5 illustrates a running example in which only the unfinished tasks are fed to the agentic system. Additionally, this algorithm is independent of the specific organization of the agentic system, as shown in Figure 1.
Figure 6: Details of our agentic system organization. (a) shows a single LLM. (b) shows the system components and their representations. The user is either a human or a self-improvement learning process. The filled colors align with the text colors in Section 4.
3.1 Filtering high-quality answers
The filter function, $\sigma$, collects earned experience $\mathcal{D}$ by filtering one answer for each newly solved task in $\mathcal{S}_{t}$ and records the success rate of this answer in $\mathcal{C}_{t}$. $\sigma$ first groups the correct answers $\mathcal{B}_{t}$ by the isomorphic abstract syntax tree (Knuth, 1968). Then, $\sigma$ groups equivalent answers using AST isomorphism and randomly selects one representative from each group. Among these representatives, $\sigma$ chooses the one with the minimal length of pure code (excluding comments and empty lines) as the final answer for the task in $\mathcal{S}_{t}$. $\sigma$ also stores the success rate from $\mathcal{C}_{t}$ for $\beta$. The minimal length selection follows the Minimum Description Length principle for higher information density (Rissanen, 1978). On the other hand, shorter text might lose chain-of-thought comments (Wei et al., 2022b). We purposefully introduce randomness to the selection of representative answers for each isomorphic group to balance these two contradicting intuitions.
3.2 Stratification and selection
The selection function $\beta$ stratifies the earned experiences $\mathcal{D}$ by binning them into $m$ levels of difficulty and demonstrations $d_{j}$ are selected incrementally from the stratified experiences $\mathcal{E}$. We define the difficulty as the opposite of the success rate following (Lightman et al., 2023). $\beta$ sorts $\mathcal{D}$ in ascending order of success rate and then bins the tasks as evenly as possible to get the boundaries of each bin. Then tasks are rebinned using these boundary values. This selection strategy can cause repetitive steps as exemplified by the dashed-line circles (Iter 4&5 in Figure 7(a)) when the newly finished tasks are easier than the demonstrations. Our algorithm keeps these repetitive steps instead of avoiding them because the tasks with low success rates have a better chance of getting one correct answer with the number of samples doubled. For example, Iter 4 performs better than Iter 3 in Figure 7(b) with the same demonstrations. This method can also cause later iterations to have fewer tokens but with higher quality than previous iterations (Iter 2&3 in Figure 7(a)) when the boundary value crosses two bins and the newly finished tasks are easier.
3.3 Discussion
Algorithm 1 can also extend to new tasks that typically involve new ML operators and new hardware specialized functions. These can be incorporated into the initial task pool $\mathcal{X}$ and handled using the same adaptive self-improvement learning process by selectively sampling only the new tasks. Each example takes about 500 tokens, so cutting-edge LLMs can handle hundreds more tasks. If the number exceeds the context length, better stratification and selection functions are needed to preserve experience quality within the context window limit. If the task set $\mathcal{X}$ is finite, as in our case, the algorithm will terminate.
4 Agentic system organization
In this section, we introduce a specific agentic system organization tailored for ML library development using STeP as shown in Figure 6. The agentic system comprises LLM agents, a code generator, and verifiers, with a structural intermediate representation as the interface between users and system components.
4.1 LLM agents
As shown in Figure 6(a), each single LLM is designed for a specific purpose assigned by the domain expert through the base prompt. The base prompt contains the task description, base knowledge, and demonstrations. The prompt composer chains the base prompt and task-specific input, which is then fed into a configured LLM serving API.
Specifically, we design two agents, a proposer and a guardian. As explained in Section 2.3, the affine type constraint in STeP is a global property that requires thinking back and forth beyond step-by-step reasoning, which is challenging for the causal generation of LLMs. Therefore, we design a guardian agent to check and correct the affine type error globally. We exclude the demonstration tasks in the base prompt of the agents from the benchmark to avoid the model directly copying the answer.
The proposer agent generates a candidate STeP implementation whose base prompt comes from ASPL designers. The base prompt is composed of STeP references and usage patterns. Figure 14 shows the reference for the Accum primitive, and Figure 17 exemplifies one of the usage patterns for shape manipulation.
The guardian agent decides whether the output of the proposer violates the affine type constraint and corrects the implementation when necessary. The guardian agent consists of a fused gate and corrector. Figure 20 shows the base prompt for the guardian agent, which provides input and output examples of variable reuse where the variable is reused zero, one, or two times.
4.2 Code generator
After the LLM agents, a code generator takes in the generated implementation and outputs a self-contained pytestable Python script. With this code generator, the LLM agents only need to output implementations without other helper code. We embed the STeP specifications written in natural language to Python (see Figure 21 and Figure 22). Beneath this Python frontend is a functional simulator that calculates the result of STeP programs. Since the essence of each ASPL is the programming abstraction (semantics) instead of its syntax (Liskov, 2011), we choose to prompt LLMs with their familiar Python syntax. In this way, LLMs can focus more on reasoning about the STeP programming abstraction without being distracted by less familiar syntax.
4.3 Verifier
Fast verification is vital because it bottlenecks adaptive self-improvement learning. ASPLs further increase this complexity by requiring a simulator for the library–chip co-design process. Simulating STeP as a general dataflow system in Python would be slow because of its high dynamism (Zhang et al., 2024b). Since we only focus on functional correctness in this work, we meticulously limit the level of dynamism to the degree that ML operators require. Consequently, our simulator emulates stream execution using tensor computation with the necessary control flow.
Our system organization contains two verifiers, and the reward function for our system in Algorithm 1 is $r(x,y)=1$ for task $y$ if answer $x$ passes these two verifiers and $r(x,y)=0$ otherwise. One verifier checks for functional correctness. Users program PyTorch to express their ML operators, which elicits LLMs’ base knowledge of ML algorithms. The verifier compares the execution results of our simulator with the result tensors of the corresponding PyTorch program on a single set of shapes with random input values. The fidelity of our unit test method builds on practice (Jia et al., 2019) and theory (Gulwani & Necula, 2003). The other verifier checks the affine type constraint by performing static analysis on the abstraction syntax tree of the STeP program with the Python ast module (Ronacher, 2008).
4.4 Structual intermediate representation
Good interfaces can improve the performance of agentic coding systems (Yang et al., 2024; Wei et al., 2024). Therefore, we are inspired by the intermediate representation (IR) technique from compiler literature (Lattner et al., 2021; Vasilache et al., 2022) and use a structural IR to unify the interfaces of our agentic system. Specifically, this structural IR is the interface between users and the agentic system and between LLM agents and the code generator within the system.
Our structural IR encodes necessary information using a data serialization language. It externalizes and condenses programs instead of simply formatting the prompt without changing the content. Comparing the structural IR in Figure 18 with the equivalent bare Python in Figure 19 for the same task, users only need to state two things: the ML operator to implement and the specialized functions as in Figure 15 without any redundant glue string. Moreover, structural IR saves tokens by reducing redundant prompts, allowing for more demonstrations.
5 Benchmark
We construct a set of tasks to measure the adaptive self-improvement agentic system proposed in Section 3 and Section 4. This benchmark should cover a diverse set of popular ML operators and specialized functions. In total, we collect 26 tasks covering 8 groups of ML operators in common LLM model architectures, as shown in Figure 3.
5.1 Metric
We choose pass@k (Chen et al., 2021) as the metric for task completion. Pass@k is calculated as Equation 1 given $T$ tasks, $n$ samples of the agentic system, and $c_{i}$ correct responses for each task $i$. It is useful for us to analyze the metric at two extremes: pass@1 and pass@n. Pass@1 is the expectation of the success rate across tasks. Pass@n is the expectation of the portion of tasks that can be solved given all samples.
$\text{pass@k}\coloneqq\frac{1}{T}\sum_{i=1}^{T}\left[1-\frac{\binom{n-c_{i}}{k}}{\binom{n}{k}}\right$] (1)
5.2 Benchmark Construction
We construct the benchmark from first principles and do not favor any kind of task. Firstly, the number of tasks in each group is nearly the same as shown in Figure 3. Secondly, the benchmark has an even distribution of difficulties. Table 2 shows that both tasks requiring shape and arithmetic primitives and tasks with and without reused variables distribute fairly evenly.
We provide a reference implementation for each task. These oracle implementations ensure that each task has at least one correct answer. Each task of one type has either different specialized functions for the same operator or different operators with different specialized functions. More details on the benchmark are in Section A.2.
6 Experiments
Detailed experimental settings are in Section A.3. We also benchmark the tasks with OpenAI-o1 in Table 4 but do not include it in the following experiments to control for test-time compute. As described in Section 4.4, all prompts are formatted in YAML because structural prompts generally benefit (He et al., 2024).
6.1 Analysis of adaptive self-improvement learning
This section uses the agentic system organization described in Section 4.1 for the best possible base learning capability. Our evaluation provides the following insights:
The hard-only examples improve performance more effectively than examples mixed with easier ones. As shown in Figure 7, the Pareto optimal is composed of hard examples (denoted by “H”) for all three models. Moreover, hard examples bring the most significant improvement along the learning curve. In some cases, fewer hard examples may perform better than more examples mixed with easier examples. For example, Iter 3 has better performance than Iter 2 for gpt-4o.
Mixed examples are required to generate better hard-only examples. In DeepSeek-V3, although HM1 performs the same as H1 and HME1 performs worse than H2 while taking more tokens, H2 would not be discovered without HM1 and HME1. Although the new hard-only examples do not necessarily improve the performance, they can save input tokens, as exemplified by HM4&H5 of gpt-4o and HM5&H6 of llama.
[…]
Adaptive granularity affects token efficiency and peak performance. If the adaptive granularity $m$ is too small, then easy examples might dilute the difficulty of training data. If $m$ is too large, then exploration steps might be too conservative and thus waste input tokens. Therefore, there is a sweet spot that balances the training data difficulty and cost of input tokens. As shown in Figure 8, $m=3$ is that spot. $m=3$ saves $1.07\times$ tokens over $m=4$ while maintaining performance. Additionally, $m=3$ improves the performance by $1.5\times$ at a similar input token cost when compared to $m=1$.
Figure 8: Hyperparameter tuning of adaptive granularity $m$ on GPT-4o. This supports using m=3 for Figure 7.
6.2 Ablation study of agentic system organization
The experiments below study the base learning capability of agentic systems without the self-improvement process.
Method Success Failure Overall Pass@n
diversity diversity diversity
Single- 0.32 0.47 0.41 0.46
w/o-IR
Single 0.27 0.64 0.50 0.62
Agent 0.34 0.68 0.52 0.85
Table 1: Analysis of the correlation between semantic diversity of answers and the performance. “-w/o-IR” means “without structural IR”. Higher values indicate higher semantic diversity.
The agentic system can discover non-trivial STeP programs. Surprisingly, the agentic system composes attention operators with over 50% Pass@1 as shown in Figure 13. That means the agentic system can discover online softmax (Milakov & Gimelshein, 2018) and memory-free streaming attention (Sohn et al., 2024b) with specialized functions provided in Section A.7, which is considered challenging for ordinary programmers.
Mode Metric Method Once Reuse Avg
Arith Pass@1 Single 0.685 0.232 0.444
Agent 0.663 0.455 0.552
Pass@n Single 7/7 5/8 12/15
Agent 7/7 8/8 15/15
Shape Pass@1 Single 0.145 0.004 0.094
Agent 0.167 0.051 0.125
Pass@n Single 3/7 1/4 4/11
Agent 4/7 3/4 7/11
Avg Pass@1 Single 0.415 0.156 0.296
Agent 0.415 0.320 0.371
Pass@n Single 10/14 6/12 16/26
Agent 11/14 11/12 22/26
Table 2: Analysis of improvement brought by the agentic method. “Arith” and “Shape” mean the oracle implementation only involves arithmetic primitives and involves shape manipulation primitives as introduced in Section 2.3, respectively. “Once” and “Reuse” mean all the streams are used once and more than once, respectively.
Our structual IR improves performance by increasing the sample diversity. Figure 9 shows the efficacy of our structural IR and agentic method. We calculate the semantic diversity of sampled answers. A group of answers is considered to have the same semantics if their abstract syntax trees are isomorphic. We calculate success diversity as the number of semantically different correct answers divided by the total number of correct answers, averaged across all tasks. Failure diversity is calculated in a similar way but for wrong answers. Overall, diversity combines both metrics. As shown in Table 1, Pass@n has a positive correlation with failure, overall diversity, and the complexity of the methods. However, structural IR can hurt success diversity.
The guardian agent can correct affine type errors but might corrupt the correct answers output by the proposer agent.
Figure 9: Pass@k against the number of samples for a single LLM without structural IR, baseline single LLM, and agentic system.
As shown in Table 2, the guardian agent effectively corrects the proposer agent, solving 5 extra reuse tasks (Pass@n of Avg-Reuse from 6/12 to 11/12). Notably, all the Arith-Reuse tasks can be solved by the agentic system (Pass@n of Arith-Reuse from 5/8 to 8/8). Surprisingly, the agentic system also helps with non-reuse tasks (Pass@n of Shape-Once from 3/7 to 4/7). However, the agentic system reduces the Pass@1 of Arith-Once from 0.685 to 0.663, implying that the guardian agent might corrupt the proposer’s output. The agentic system can compensate for such corruption by finishing more tasks, resulting in an unchanged Pass@1 of Avg-Once.
7 Related Work
This work introduces a self-improvement agentic system for ASPL code generation, enhancing LLM effectiveness through an adaptive learning algorithm and deliberate agentic system design. Accordingly, we focus the related work on approaches that leverage LLMs for self-improvement and agentic systems on specialized tasks.
7.1 Self-improvement learning for LLMs
Self-improvement learning for LLMs typically involves two stages: scoring generated samples (trajectories) and incorporating those samples to enhance the model. Scoring can be achieved through human labeling (Cobbe et al., 2021; Lightman et al., 2023) or through automated methods such as verifiers and heuristics (Wang et al., 2024; Singh et al., 2023). Our method stands out in this context by utilizing AST analysis, offering a more interpretable approach to scoring. When it comes to incorporating samples, models may rely on retrieval (Zhao et al., 2024; Park et al., 2023), reflection (Shinn et al., 2024; Liu et al., 2023; Yao et al., 2023), or reward feedback (Opsahl-Ong et al., 2024; Fernando et al., 2023). Our method introduces a new mechanism in this stage by adaptively extending and prioritizing high-scoring samples.
Our method shares similar reinforcement principles with self-improvement learning at the post-training stage (Bai et al., 2022; Gulcehre et al., 2023; Tian et al., 2024) but is rewarded at a task-level granularity instead of token-level. Specifically, each action is a program implementation instead of token prediction and the state is defined by earned experiences rather than generated sequences.
7.2 Agentic system organization for specialized tasks
Task-specific organization has proven effective in enhancing the performance of agentic systems across diverse coding tasks (Zhang et al., 2024c; Fang et al., 2024; Guan et al., 2024). We adopt an agentic system organization specifically for ML library development using an ASPL. Such domain knowledge can be further augmented with automatic agentic system design tools (Khattab et al., 2023; Hu et al., 2024; Zhang et al., 2024a). Furthermore, well-designed interfaces between agents, tools, and other agents have been shown to improve performance (Schick et al., 2023; Yang et al., 2024; Wu et al., 2023), and a structural IR enables these interfaces to be highly task-aligned in our system.
8 Conclusion
ML library development using ASPLs is a critical component of the ML ecosystem, but it remains poorly automated. To address this limitation, we co-design the learning process and agentic system around a central objective: enabling complex reasoning with limited data. Our methods simultaneously implement non-trivial ML operators and produce a self-improving agent. Consequently, our system not only supports experts in developing ML libraries but also offers valuable resources for other systems. We recognize that the challenges of complex reasoning under limited data extend beyond this domain and envision applying similar self-improvement techniques with LLM agents to a broad class of tasks.
Impact Statement
Our work on adaptive self-improvement learning LLM agentic systems for automating ML library development using ASPLs has several potential societal implications. The primary impact is the potential to significantly enhance the productivity of ML library developers, which could accelerate the development of more efficient ML systems. This advancement could democratize access to ML development tools and reduce the technical barriers to entry in the field. Additionally, our technique offers an approach for deploying LLM agents in scenarios requiring complex reasoning with limited data availability. While these developments primarily aim to advance the field of Machine Learning, we acknowledge that increased automation in software development could impact the nature of programming work and skills required in the field. We believe these potential implications warrant ongoing discussion and careful consideration as the technology develops.
Acknowledgement
We thank Vishnu Sarukkai, Haochen Shi, Xinhao Li, Gina Sohn, Nathan Zhang, Tian Zhao, Yu Sun, Anjiang Wei, Chun Deng, Qizheng Zhang, Rubens Lacouture, Fredrik Kjolstad and the anonymous reviewers for their feedback on this paper. This research was supported in part by the Pervasive Parallelism Lab Affiliates Program Fund. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the aforementioned funding agencies.
Appendix A Appendix
A.1 STeP introduction
$\sum_{i=0}^{1}G_{i}\cdot\text{gelu}(W_{i}X)\text{ with }N_{i}=\mathbf{I}[G_{i}>0]$ (2)
Algorithm 2 STeP for a simplified MoE module
0:$X$: [m,n] of Buffer(k), NN: [m,n] of Multihot(e), $G$: [m,n] of Buffer(e)
0: [m,n] of Buffer(k)
[…]
0: weightedsum $\triangleright$ external function
[…]
1:fn (s,v) = s + $v_{0}*v_{1}$
1:init: s = 0
0: expert0
[…]
1:fn (v) = gelu($W_{0}v_{0}$), $v_{1}$[0])
0: expert1
[…]
1:fn (v) = gelu($W_{1}v_{0}$), $v_{1}$[1])
[…]
Figure 10: Algorithm 2 is a STeP program example for Equation 2. The type signature follows the Haskell style where $\rightarrow$ connects a sequence of argument types with one return type. Three funcs are external functions provided by the hardware.
As shown in Figure 10, Copy duplicates a stream for the affine type constraint. Zip combines two streams of values into a stream of tuples. Map applies the function (fn) on each input value. Partition routes tokens of the data stream to experts assigned by the index stream ($S_{1}^{0}$). Merge accumulates tokens from experts. The accumulation of Merge is parameterized by fn and init where init initializes the state and fn updates the state with the input value. Partition and Merge are used in pairs, sharing the same index stream. $\langle\rangle$ represents the Tuple type of value tokens. Buffer, Multihot, and Scalar are also types of value tokens, parametrized by the generic data type like float and half, which is omitted for simplicity in the algorithm. Buffer and Multihot types are further parametrized by the shape in the paratheses. Shape manipulation primitives include Promote, Repeat and RepeatRefFigure 4(a) also shows two streams. S1 is a control token signaling the end of rank-1. 01 is a value token of multihot vector type served in index streams. (v0,v1) is a value token of type Tuple(Scalar, Reference) because the weight and input streams are composed of scalar and reference values, respectively.
A.2 Benchmark details
As shown in Figure 3, Group 4 tasks have the same operator: $softmax(S)\cdot V$ where $S$ equals $QK^{T}$. They differ in external functions as shown in Section A.7. Group 4 can use RDA’s on-chip fusion to compose scale-dot-product attention with Group 5. Group 7 also differs in external functions. Group 5 contains three dataflow orders: inner-product($mnk,mdk\rightarrow mnd$”), row-wise($mnk,mkd\rightarrow mnd$”), and outer-product($mkn,mkd\rightarrow mnd$”). Group 6 contains GptJ and NeoX styles which differ in pairing even and odd or the first and second half positions (vLLM, 2023). Group 8 contains LayerNorm and RMSNorm.
The last three in Table 3: index, expert, and etoe all come from MoE. MoE contains token (Shazeer et al., 2017) Equation 3 and expert choice routing (Zhou et al., 2022) Equation 4.
$S=softmax(X\cdot W_{g}),S\in\mathbb{R}^{n\times e}$ (3)
$G,I=TopK(S)$ Along expert dimension
$S=softmax(X\cdot W_{g}),S\in\mathbb{R}^{n\times e}$ (4)
$G,I=TopK(S^{T})$ Along token dimension
The expert choice routing can have another MLP auxiliary predictor for causal inference when it is a binary choice  (Raposo et al., 2024) Equation 5.
$S=\sigma((gelu(X\cdot W_{g_{0}}))\cdot W_{g_{1}}),S\in\mathbb{R}^{n\times 1}$ expert=2 (5)
[…]
Group Description Count
attn Softmax(S)@V part 3
gemm Matrix multiplication with expanded reduction dimension 3
bmm Batch matrix-matrix product 3
norm RMSNorm and LayerNorm (without bias and gain) 3
rope GptJ and NeoX style of RoPE 3
index Index generation of MoE router 4
expert Expert execution of MoE 3
etoe End-to-end of MoE module 4
Table 3: Description of each type of tasks. Matmul is short for matrix multiplication.
A.3 Experiment settings
In Section 6.1 we sample 64 times for each temperature of 0.4, 0.7, and 1.0, recording the best result. In Table 1, we sample 64 times at temperature 0.7 on Claude 3.5 Sonnet to control variables. Four models are: claude-3-5-sonnet-20241022 of Anthropic API (Claude 3.5 Sonnet), gpt-4o-2024-11-20 of OpenAI API (GPT-4o), deepseek-chat of DeepSeek API (DeepSeek-V3), and Meta-Llama-3-1-405B-Instruct-Turbo of TogetherAI API (Llama 3.1-405B). Maximum output tokens are set as 1024 and the seed for GPT-4o is 42.
A.4 Additional experiments
Figure 11: Self-improvement learning curve with m=3. Claude 3.5 Sonnet consumes much fewer tokens than other models because the input tokens are counted by the least necessary number of tokens averaged across tasks. A task can take increasing input tokens but still fail. Sonnet only has 3 unsolved tasks left. Therefore, although it has the most example tokens, the average number of input tokens across tasks is still less than others.
Figure 12: Result of five models at temperature 0.7.
Figure 13: Result of three methods on Claude 3.5 Sonnet at temperature 0.7.
As shown in Table 4, OpenAI-o1 achieves similar performance at the cost of more tokens than a single Claude-3-5-Sonnet. This observation aligns with previous findings that scaling pretraining is preferable over inference for challenging tasks (Snell et al., 2024). Meanwhile, a single Claude-3-5-Sonnet proposer can finish more tasks than the OpenAI-o1 using fewer tokens. Table 3 contains abbreviations for all tasks.
Model Metric ($\uparrow$) attn gemm bmm norm rope index expert etoe avg
Qwen2.5-Coder-32B Pass@1 0.010 0 0.052 0 0 0.004 0 0 0.008
Pass@n 0.33 0 0.33 0 0 0.25 0 0 0.115
Llama3-405B Pass@1 0.010 0 0 0.057 0.089 0.016 0 0 0.020
Pass@n 0.33 0 0 1.00 0.67 0.25 0 0 0.269
DeepSeek-V3 Pass@1 0.438 0 0 0 0 0.113 0 0 0.068
Pass@n 1.00 0 0 0 0 0.75 0 0 0.231
GPT-4o Pass@1 0.021 0.016 0.214 0.094 0 0.258 0.005 0 0.080
Pass@n 0.33 0.67 0.33 0.67 0 0.5 0.33 0 0.346
Claude-3-5-sonnet Pass@1 0.620 0.229 0.146 0.208 0.526 0.676 0.688 0.324 0.433
Pass@n 1.00 1.00 0.33 1.00 1.00 0.75 1.00 1.00 0.885
OpenAI-o1 (n=8) Pass@1 0.208 0.042 0 0 0.083 0.343 0.583 0 0.159
Pass@n 0.67 0.33 0 0 0.67 0.5 1.00 0 0.385
Table 4: The performance of the self-improvement agentic system across models.
Our system completed each task in under 10 minutes on average. In contrast, during our pilot study, a domain expert was unable to write a single program within 48 hours, as they had to do trial-and-error and accumulate experience sequentially. Our system, by comparison, can perform these explorations in parallel. In the future, collaboration with HCI researchers will enable more extensive experiments comparing time and effort required by human programmers versus the system, providing quantitative data on usability and cognitive load.
Differences in tool access and computational load might have influenced the outcomes. Therefore, we conducted an experiment that aligned both aspects.
For computation fairness, we matched the token count of the single model with the agent and self-improved models by resampling. All model variants (single, agent, self-improved) have access to the same verifier so it is fair; the difference lies in how the verifier is leveraged. Self-improved models incorporate it throughout the process, while others use it only at the end as a final judge. We chose Claude-3-5-Sonnet and GPT-4o as base models. As shown in Table 5, our agentic systems still perform better under this fair setting.
Pass@n Claude Sonnet Claude Sonnet GPT-4o GPT-4o
Single Agent & Self-improved Single Agent & Self-improved
From 0.73 0.73 0.23 0.23
To 0.77 0.96 0.38 0.81
Table 5: Performance comparison of single models vs agent & self-improved models with token count matching.
We conducted supervised finetuning (SFT) using GPT-4o. Table 6 shows that SFT can improve performance, but less than our self-improvement approach.
Since we do not know the exact SFT algorithm of OpenAI service for FLOPs matching, we tried our best to favor the SFT method. We began with the same 133 correct samples from all completed tasks used in the first iteration of self-improvement. Different from self-improvement which only picks 1 correct program per completed task, we picked all 133 programs to form the training dataset of SFT. We created three SFT datasets with varying prompt compositions:
  • 133 (base prompt+question+answer)
  • 133 (question+answer)
  • 17 (question+answers deduplicated via AST)
Each dataset was used to train a separate SFT model. After that, we sampled each model on all the uncompleted tasks.
Pass@n Finetuned Self-improved
From 0.35 0.35
To 0.62 0.81
Table 6: Performance comparison between finetuned and self-improved Models
We assessed code maintainability using two metrics: maintenance index without comments (MIwoc) and with comments (MI)  [^2] https://www.verifysoft.com/en_maintainability.html. The comment weight (MIcw) is defined as MI - MIwoc and falls in [0, 49); MI >> 85 indicates good maintainability. Using all correct programs from our best model (self-improved agentic Claude Sonnet), we recorded the top MIwoc and MI per task. The mean MIwoc is 102, MI is 149, and MIcw is 47—indicating well-commented, maintainable code.
As shown in Table 7, we used the same set of programs as the maintainability statistics to measure complexity. We measured these three metrics:
  • Lines of code: Counted via primitive calls (excluding comments/blank lines)
  • Shape transformations: Counted by use of Promote, Repeat, RepeatRef, and Flatten primitives
  • Specialized instructions: Counted as the number of specialized functions in task descriptions
Metric Min Max Mean
Lines of code 4 17 8.67
Shape transformations 0 6 1.13
Specialized instructions 2 7 3.68
Table 7: The assessment of complexity across all completed tasks.
A.5 Generalization
The proposed system can be extended to other scenarios that require complex reasoning with limited example data and well-defined evaluation metrics. We outline the general recipe below.
As shown in Figure 1, the agentic system organization is constructed in three main steps. First, system designers define the format of both the task and its expected output. Once the format is specified, the next step is to build a verifier for the task. With the format and verifier in place, designers can either use a single LLM or design LLM agents tailored to the domain—similar to how we handle the type constraints of the STeP language. After completing these three steps, the task can be handed over to our system, which will automatically carry out adaptive self-improvement learning.
The adaptive self-improvement learning system also exposes several tunable hyperparameters which will be helpful when the results are not satisfactory. The most direct control is the number of parallel sampling. Users can also adjust the adaptive granularity parameter m for experience stratification. Additionally, domain-specific filtering heuristics—such as the minimal code length heuristic used by us—can be incorporated to further guide the learning process.
Our approach has two parts: adaptive self-improvement learning and agentic system organization. The learning process is broadly applicable to other programming languages; the challenges lie in tailoring agentic systems to other languages.
Mainstream languages like CUDA, HIP, and CPU vector intrinsics exhibit global properties such as arbitrary memory access, data layout sensitivity, and side effects. Similarly, STeP enforces a global affine type constraint. Our framework addressed this using a guardian agent that detects and corrects affine type violations. This concept generalizes: domain-specific guardian agents can monitor and enforce global properties of various languages, adapting the STeP solution more broadly.
A second challenge is that LLMs may lean toward surface-level patterns in mainstream languages due to their existence in training data, potentially missing more optimal or novel transformations. As shown in Section 6.2, our structural IR can increase sample diversity and thus boost the LLM agentic system performance. Extending this, structural IRs and tailored code generators can guide LLMs toward more creative solutions beyond conventional patterns.
We also conducted an experiment on the AIME-2024 dataset  [^3] https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 which contains 30 challenging problems from the American Invitational Mathematics Examination (AIME) 2024. We applied our adaptive self-improvement learning to the Claude-3-5-Sonnet base model and increased Pass@n from 0.50 to 0.67. This demonstrates the potential capabilities of our system on other tasks.
A.6 Additional explanation
Abbv. Description
DSA Domain Specific Architectures
DSL Domain Specific Language
ASPL Architecture Specific Programming Language
RDA Reconfigurable Dataflow Architecture, a DSA for AI
STeP Streaming Tensor Program, an ASPL for next-generation RDA
Table 8: Explanation of the abbreviations.
Existing auto-tuning code generation systems like Spiral (Puschel et al., 2005), Ansor (Zheng et al., 2020), and AKG (Zhao et al., 2021) rely on predefined optimization strategies and cost models. On the contrary, we employ LLMs to generate and refine code for machine learning libraries, particularly targeting ASPLs. We explore the shift towards utilizing the reasoning capabilities of LLMs for code generation and optimization, enabling the system to adapt and improve over time without explicit human intervention.
A.7 Code and prompt examples
- name: Accum
desc: |
Accum is a primitive operation that applies a function to a stream in a recursive manner.
The function is applied to the first element of the stream and the initial state to produce the first output element.
The function is then applied to the second element of the stream and the output of the previous application to produce the second output element, and so on.
The state is initialized at rank b of the input stream. The output streams shape is the input streams shape excluding the first b dimensions.
examples:
- inputs:
- name: E0
dtype: fp32
dims: [M, N]
data_gen: torch.rand
fns:
- name: Sum
apply: |
return [state[0] + input[0]]
init: [0]
input_dtype: fp32
output_dtype: fp32
func_name: fn_sum
outputs:
- name: S0
dtype: fp32
dims: [N]
data_transform:
torch.sum(input_data[’E0’], 1, keepdim=False)
impl: |
E1 = step.Accum(fn=fn_sum, b=1).apply(E0)
return E1
Figure 14: The reference of Accum that contains the definition and an example. Each example in the examples field is composed of task description and implementation.
inputs:
- name: E0
dtype: fp32
dims: [M, N]
data_gen: torch.rand
- name: E1
dtype: Buffer(fp32, [D])
dims: [M, N]
data_gen: torch.rand
fns:
- name: MaxSum
apply: |
m_t, l_t, o_t = state # scalar, scalar, [D]
s_t, v_t = input # scalar, [D]
m_next = torch.max(m_t, s_t) # scalar
l_prim_t = torch.exp(m_t - m_next) * l_t
p_t = torch.exp(s_t - m_next)
l_next = p_t + l_prim_t
o_next = l_prim_t * o_t / l_next + p_t * v_t / l_next
return [m_next, l_next, o_next]
init: [-inf, 0, 0]
input_dtype: [fp32, "Buffer(fp32, [D])"]
output_dtype: [fp32, fp32, "Buffer(fp32, [D])"]
func_name: fn_maxsum
- name: GetThird
apply: |
return [input[2]]
input_dtype: [fp32, fp32, "Buffer(fp32, [D])"]
output_dtype: Buffer(fp32, [D])
func_name: fn_getthird
outputs:
- name: S0
dtype: fp32
dims: [D, N]
data_transform:
torch.bmm(torch.softmax(input_data[’E0’], 1).unsqueeze(1), input_data[’E1’]).squeeze(1)
impl: |
Figure 15: Task description of attn task 0 in our structural IR, where the LLM needs to complete impl.
# Task 0 # MaxSum def apply(state, input): $m_{t},l_{t},\boldsymbol{o_{t}}$ = state $s_{t},\boldsymbol{v_{t}}$ = input $m_{t+1}=max(m_{t},s_{t})$ $l^{\prime}_{t}=l_{t}*e^{(m_{t}-m_{t+1})}$ $p_{t}=e^{(s_{t}-m_{t+1})}$ $l_{t+1}=p_{t}+l^{\prime}_{t}$ $\boldsymbol{o_{t+1}}=\frac{1}{l_{t+1}}(l^{\prime}_{t}*\boldsymbol{o_{t}}+p_{t}*\boldsymbol{v_{t}})$ return $(m_{t+1},l_{t+1},\boldsymbol{o_{t+1}})$ def init(): return $(-\infty,0,\boldsymbol{0})$ # GetThrid def apply(input): return input[2] # Task 1 # ExpMaxDiff def apply(state, input): $m_{t},e_{t},d_{t}$ = state $s_{t},$ = input $m_{t+1}=max(m_{t},s_{t})$ $\Delta m=m_{t}-m_{t+1}$ $e_{t+1}=e^{(s_{t}-m_{t+1})}$ $d_{t+1}=e^{\Delta m}$ return $(m_{t+1},e_{t+1},d_{t+1})$ def init(): return $(-\infty,0,0)$ # DivSum def apply(state, input): $\boldsymbol{v_{t}},e_{t},d_{t}$ = input $l_{t},\boldsymbol{o_{t}}$ = state $l^{\prime}_{t}=d_{t}*l_{t}$ $l_{t+1}=e_{t}+l^{\prime}_{t}$ $o_{t+1}=\frac{1}{l_{t+1}}(l^{\prime}_{t}*\boldsymbol{o_{t}}+e_{t}*\boldsymbol{v_{t}})$ return $(l_{t+1},o_{t+1})$ def init(): return $(0,\boldsymbol{0})$ # GetSecondThrid def apply(input): return input[1],input[2] # GetSecond def apply(input): return input[1] # Task 2 # ExpMaxDiff def apply(state, input): $m_{t},e_{t},d_{t}$ = state $s_{t},$ = input $m_{t+1}=max(m_{t},s_{t})$ $\Delta m=m_{t}-m_{t+1}$ $e_{t+1}=e^{(s_{t}-m_{t+1})}$ $d_{t+1}=e^{\Delta m}$ return $(m_{t+1},e_{t+1},d_{t+1})$ def init(): return $(-\infty,0,0)$ # GetSecondThrid def apply(input): return input[1],input[2] # WeightedSumSingle def apply(state, input): $e_{t},d_{t}$ = input $r_{t}$ = state return $(r_{t}*d_{t}+e_{t})$ def init(): return 0 # WeightedSumDouble def apply(state, input): $\boldsymbol{v_{t}},e_{t},d_{t}$ = input return $(state*d_{t}+e_{t}*\boldsymbol{v_{t}})$ def init(): return $\boldsymbol{0}$ # Div def apply(input): $r_{t},\boldsymbol{l_{t}}=input$ return $\frac{\boldsymbol{l_{t}}}{r_{t}}$
Figure 16: Inner functions for 3 tasks of attn. Task 0 encapsulates the whole innermost loop body of FlashAttention (Dao et al., 2022) in the MaxSum function. Task 1 splits the MaxSum into ExpMaxDiff and DivSum. Task 2 postpones the division of summation as in FlashAttention2 (Dao, 2023). The bold symbols are streams with type 1D Buffer, and the symbols are streams with type Scalar.
name: Stashing dimension
desc: |
When the pritmives require a non-one dimension to be inserted as a non-innermost dimension, a Bufferize&Streamify pair can wrap the primitives to adjust the dimension.
This pattern is useful for Repeat and RepeatRef primitives.
examples:
- inputs:
- name: E0
dtype: fp32
dims: [M, N, K]
data_gen: torch.rand
outputs:
- name: S0
dtype: fp32
dims: [M, N, D, K]
data_transform:
input_data[’E0’].unsqueeze(1).repeat(1, D_value, 1, 1)
impl: |
E1 = step.Bufferize(a=2).apply(E0) # E1: {dtype: Buffer(fp32, [M, N]), dims: [K]}
E2 = step.Repeat(n=D).apply(E1) # E2: {dtype: Buffer(fp32, [M, N]), dims: [D, K]}
E3 = step.Streamify().apply(E2) # E3: {dtype: fp32, dims: [M, N, D, K]}
return E3
Figure 17: An example of usage pattern that contains 3 shape manipulation primitives: Bufferize, Repeat, and Streamify.
inputs:
- name: E0
dtype: fp32
dims: [M, N]
data_gen: torch.rand
- name: E1
dtype: Buffer(fp32, [D])
dims: [M, N]
data_gen: torch.rand
fns:
- name: MaxSum
apply: |
m_t, l_t, o_t = state # scalar, scalar, [D]
s_t, v_t = input # scalar, [D]
m_next = torch.max(m_t, s_t) # scalar
l_prim_t = torch.exp(m_t - m_next) * l_t
p_t = torch.exp(s_t - m_next)
l_next = p_t + l_prim_t
o_next = l_prim_t * o_t / l_next + p_t * v_t / l_next
return [m_next, l_next, o_next]
init: [-inf, 0, 0]
input_dtype: [fp32, "Buffer(fp32, [D])"]
output_dtype: [fp32, fp32, "Buffer(fp32, [D])"]
func_name: fn_maxsum
- name: GetThird
apply: |
return [input[2]]
input_dtype: [fp32, fp32, "Buffer(fp32, [D])"]
output_dtype: Buffer(fp32, [D])
func_name: fn_getthird
outputs:
- name: S0
dtype: fp32
dims: [D, N]
data_transform:
torch.bmm(torch.softmax(input_data[’E0’], 1).unsqueeze(1), input_data[’E1’]).squeeze(1)
impl: |
E3 = step.Zip().apply((E0, E1))
E4 = step.Accum(fn=fn_maxsum, b=1).apply(E3)
E5 = step.Map(fn=fn_getthird).apply(E4)
E2 = step.Streamify().apply(E5)
return E2
Figure 18: The complete implementation of attn task 0 written in structural IR.
import step
from sympy import Symbol
import torch
M = Symbol("M")
N = Symbol("N")
K = Symbol("K")
D = Symbol("D")
M_value = 5
N_value = 7
K_value = 9
D_value = 11
ctx = {M: M_value, N: N_value, K: K_value, D: D_value}
input_dtype = {’E0’: step.Scalar("float"), E1’: step.Buffer(step.Scalar("float"), [D])}
input_data = {’E0’: torch.rand(N_value, M_value), E1’: torch.rand(N_value, M_value, D_value)}
class MaxSum(step.Fn):
def __init__(self, input, output):
super().__init__("MaxSum", input, output)
def getInit(self):
return [torch.tensor(float(’-inf’)), torch.tensor(0), torch.zeros(D_value)]
def apply(self, state, input):
m_t, l_t, o_t = state # scalar, scalar, [D]
s_t, v_t = input # scalar, [D]
m_next = torch.max(m_t, s_t) # scalar
l_prim_t = torch.exp(m_t - m_next) * l_t
p_t = torch.exp(s_t - m_next)
l_next = p_t + l_prim_t
o_next = l_prim_t * o_t / l_next + p_t * v_t / l_next
return [m_next, l_next, o_next]
fn_maxsum = MaxSum(step.STuple((step.Scalar("float"), step.Buffer(step.Scalar("float"), [D]))), step.STuple((step.Scalar("float"), step.Scalar("float"), step.Buffer(step.Scalar("float"), [D]))))
class GetThird(step.Fn):
def __init__(self, input, output):
super().__init__("GetThird", input, output)
def apply(self, input):
return [input[2]]
fn_getthird = GetThird(step.STuple((step.Scalar("float"), step.Scalar("float"), step.Buffer(step.Scalar("float"), [D]))), step.Buffer(step.Scalar("float"), [D]))
def prepare():
E0 = step.Stream("E0", step.Scalar("float"), 1, [M, N])
E0.ctx = ctx
E0.data = [input_data[’E0’]]
E1 = step.Stream("E1", step.Buffer(step.Scalar("float"), [D]), 1, [M, N])
E1.ctx = ctx
E1.data = [input_data[’E1’]]
return E0, E1
def check_shape(S0):
assert S0.shape == [D, N]
assert S0.dtype == step.Scalar("float")
def check_data(S0):
S0_data_0 = torch.bmm(torch.softmax(input_data[’E0’], 1).unsqueeze(1), input_data[’E1’]).squeeze(1)
torch.testing.assert_close(S0.data[0], S0_data_0)
def test():
E0, E1 = prepare()
S0 = body(E0, E1)
check_shape(S0)
check_data(S0)
def body(E0, E1):
E3 = step.Zip().apply((E0, E1))
E4 = step.Accum(fn=fn_maxsum, b=1).apply(E3)
E5 = step.Map(fn=fn_getthird).apply(E4)
E2 = step.Streamify().apply(E5)
return E2
Figure 19: The unit test of the implementation of attn task 0 in Python produced by the code generator from structural IR shown in Figure 18.
desc: |
Streaming Tensor Programs (STeP) provides a higher-level abstraction for dataflow systems.
The streams can be only consumed once. Your task is to use Copy primitives to create a new stream that is a copy of the input stream when necessary.
examples:
- input_impl: |
E2 = step.Partition(N=E_value).apply((E0, E1))
E3 = [step.Map(fn=fn).apply(s) for fn, s in zip(matmul_fns, E2)]
E4 = step.Merge(fn=fn_sum).apply((E3, E1))
return E4
output_impl: |
E1_0, E1_1 = step.Copy().apply(E1)
E2 = step.Partition(N=E_value).apply((E0, E1_0))
E3 = [step.Map(fn=fn).apply(s) for fn, s in zip(matmul_fns, E2)]
E4 = step.Merge(fn=fn_sum).apply((E3, E1_1))
return E4
explanation: |
Stream E1 is consumed twice in the input implementation. To ensure that the stream is consumed only once, we create a copy of the stream E1 and use the copy in the second step.
- input_impl: |
E1 = step.Map(fn=fn_predict).apply(E0)
E2 = step.Map(fn=fn_router).apply(E1)
E3 = step.Map(fn=fn_affinity).apply(E0)
E4 = step.Zip().apply((E0, E3))
return E4
output_impl: |
E0_0, E0_1 = step.Copy().apply(E0)
E0_2, E0_3 = step.Copy().apply(E0_0)
E1 = step.Map(fn=fn_predict).apply(E0_1)
E2 = step.Map(fn=fn_router).apply(E1)
E3 = step.Map(fn=fn_affinity).apply(E0_2)
E4 = step.Zip().apply((E0_3, E3))
return E4
explanation: |
Stream E0 is consumed 3 times in the input implementation. To ensure that all streams are consumed only once, we create a copy of the stream E0 and use the copy in the subsequent steps.
- input_impl: |
E1 = step.Bufferize(a=1).apply(E0)
E2 = step.Map(fn=fn_gate).apply(E1)
E3 = step.Map(fn=fn_top2).apply(E2)
return E3, E2
output_impl: |
E1 = step.Bufferize(a=1).apply(E0)
E2 = step.Map(fn=fn_gate).apply(E1)
E3 = step.Map(fn=fn_top2).apply(E2)
return E3, E2
explanation: |
All streams are consumed only once in the input implementation. No need to create a copy of any stream.
Figure 20: Base prompt for the guardian agent.
### <a name="Accum"></a>Accum
Accumulate the lower b dimensions in Stream<A,a>‘ into a single value of type B‘. **Accum** will continue to dequeue and accumulate to a value of type B by calling the given accumulation function (‘Fn(A,B)->B‘) until it sees a ‘.Sb in the input stream. Then, it will emit the accumulated value of type B into the output stream and initialize the accumulator with the given initialize function.
Accum<A,B,a,b>: Fn(A,B) -> B, Fn() -> B, Stream<A,a> -> Stream<B,a-b>
(accumulate) (initialize)
Precondition: 0 < b <= a
We can think of b as the minimum stop token level **Accum** has to see before emitting the accumulated values. More details on how to set b according to the type of reduction we do can be found in the below examples.
<details>
<summary>
Examples
</summary>
**Example1: Rowmax** <br/>
Goal: [B,N,E] -> [B,N] (Reduce over the inner-most dim)
Accum<A=f32, B=f32, a=3, b=1>:
Fn(f32,f32)->f32, Fn()->f32, Stream<f32,3>->Stream<f32,2>
Precondition: 0 < b <= a
(=1) (=3)
We will call the given function (max) on every dequeue and emit the accumulated value when we see a ‘.Sx(x>=b)‘. b is 1 in this example because we have to see the whole vector to obtain the reduced value.
<br/>
Figure 21: The specification of Accum primitive in the STeP document.
class Accum(OpBase):
def __init__(self, **kwargs):
super().__init__("Accum", **kwargs)
def apply(self, input: base.Stream, name=""):
b = self.config["b"]
fn: base.Fn = self.config["fn"]
assert isinstance(fn, base.Fn), f"Accum should take one of provided fns as input, but get {type(fn)}"
assert fn.input == input.dtype, f"Accum should take {fn.input} as input, but get {input.dtype}"
assert b > 0 and b <= input.rank, f"Accum should take a positive integer b less than or equal to the rank of the input, but get b: {b} and input rank: {input.rank}"
result = base.Stream(
self.getName(name), fn.output, input.rank - b, input.shape[b:]
if input.data is not None:
result.ctx = input.ctx
# TODO: Construct a general application function here
output_indices = get_full_indices(base.subsOuterShape(result.shape, result.ctx))
input_indices = get_full_indices(base.subsOuterShape(input.shape[:b], input.ctx))
if isinstance(result.dtype, base.Element) or isinstance(result.dtype, base.Buffer):
output_shapes = [base.subsFullShape(result.dtype, result.shape, result.ctx)[::-1]]
elif isinstance(result.dtype, base.STuple):
output_shapes = [base.subsFullShape(r, result.shape, result.ctx)[::-1] for r in result.dtype]
else:
raise ValueError("Invalid dtype")
result.data = [torch.zeros(shape) for shape in output_shapes]
for idx in output_indices:
state = fn.getInit()
for i in input_indices:
full_idx = idx + i
partial_data = [d[full_idx + (...,)] for d in input.data]
state = fn.apply(state, partial_data)
for n, s in enumerate(state):
result.data[n][idx] = s
return result
Figure 22: The definition of Accum primitive in the Python frontend.